OpenStudy (anonymous):

What is the apothem and perimeter of a decagon inside a circle if the area is 25 times pi ?

OpenStudy (zepp):

owait, nvm :s Find the radius of the circle first :D

OpenStudy (zepp):

http://puu.sh/q9Ii There, once you found the radius, you get this, find the angle between the two r's, (hint: decagon, circle, 360*) then use the cos' law to find the base's measure, and then... and then.. you've got everything to solve this, good luck! ;D

OpenStudy (anonymous):

Thanks! didnt give me a radius though!

OpenStudy (zepp):

Aha! You must find the radius by yourself! ^_^ What's the formula to find a circle's radius?

OpenStudy (anonymous):

No clue... :/

OpenStudy (zepp):

A circle's radius is pi * r^2

OpenStudy (zepp):

We know that the area is 25pi So set pi * r^2 equal to 25pi 25 pi = pi r*2 Solve it, and you'll find the radius :)

OpenStudy (zepp):

25 pi = pi r*2 25 = r^2 r = sqrt(25) r= 5 For the angle, since it's a decagon, you divide 360 by 10, making 36 degrees. Now you have an angle and two measures by its side c^2 = a^2 + b^2 - 2ab cos(C) By using the law of cosine, the base measures ~7.5

OpenStudy (anonymous):

Thanks so much!

OpenStudy (zepp):

Now you have the base, find the apothem using the pythagorean theorem, good luck ;D

OpenStudy (anonymous):

Thanks. :)